# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v3

@@ H.4 More Sensitivity Analysis
WL subtree 71.6 72.1 62.1 71.2 57.8 67.7 64.0 75.3 41.1 59.2 55.9 55.4 62.8
GCN 66.5$\pm$0.4 60.0$\pm$0.8 55.4$\pm$0.3 60.0$\pm$0.1 39.6$\pm$0.3 40.0$\pm$0.4 55.4$\pm$0.2 66.5$\pm$0.1 39.6$\pm$0.6 33.5$\pm$0.3 39.6$\pm$0.1 44.7$\pm$0.2 50.1
GIN 71.4$\pm$4.7 73.4$\pm$3.4 60.8$\pm$2.7 66.0$\pm$3.4 50.5$\pm$3.7 51.6$\pm$1.8 64.8$\pm$1.0 71.3$\pm$3.5 48.3$\pm$4.2 57.4$\pm$3.8 55.1$\pm$3.4 52.6$\pm$4.3 60.3
GMT 67.4$\pm$1.0 61.7$\pm$2.1 55.8$\pm$0.7 57.0$\pm$2.4 60.2$\pm$0.5 58.2$\pm$2.0 57.8$\pm$2.1 65.7$\pm$1.3 60.2$\pm$0.3 57.3$\pm$2.3 60.7$\pm$0.6 57.1$\pm$1.2 59.9
CIN 70.4$\pm$2.8 66.5$\pm$4.3 58.5$\pm$2.6 64.2$\pm$2.7 60.6$\pm$3.0 64.2$\pm$3.2 58.7$\pm$2.4 69.1$\pm$2.7 57.5$\pm$3.4 67.7$\pm$2.1 59.5$\pm$2.3 56.1$\pm$1.2 62.7
SpikeGCN 66.5$\pm$0.9 60.1$\pm$1.5 55.9$\pm$0.8 60.2$\pm$0.6 54.8$\pm$2.1 59.9$\pm$1.6 55.8$\pm$0.8 62.9$\pm$1.6 58.4$\pm$1.2 61.1$\pm$1.3 58.8$\pm$1.9 62.1$\pm$1.5 59.8
DRSGNN 67.5$\pm$1.2 61.2$\pm$0.7 55.6$\pm$1.4 61.1$\pm$0.9 52.4$\pm$2.3 61.0$\pm$1.2 56.9$\pm$0.7 66.7$\pm$1.5 60.3$\pm$0.5 62.0$\pm$2.0 59.8$\pm$1.4 59.2$\pm$1.1 60.3
CDAN 72.9$\pm$0.2 74.0$\pm$0.3 62.7$\pm$0.3 73.8$\pm$0.5 61.2$\pm$1.0 70.0$\pm$1.2 62.8$\pm$0.1 73.0$\pm$0.3 60.6$\pm$0.2 71.6$\pm$1.5 60.5$\pm$0.2 61.1$\pm$1.4 67.0
ToAlign 68.0$\pm$3.8 73.4$\pm$2.7 64.5$\pm$1.1 63.7$\pm$2.4 60.6$\pm$1.2 61.9$\pm$1.3 64.8$\pm$1.3 74.0$\pm$1.3 60.0$\pm$0.6 65.7$\pm$3.1 61.0$\pm$1.4 56.2$\pm$2.3 64.5
MetaAlign 73.6$\pm$0.2 72.7$\pm$1.9 63.9$\pm$1.0 67.9$\pm$4.3 60.4$\pm$0.7 65.4$\pm$1.8 65.2$\pm$0.8 73.2$\pm$2.3 60.0$\pm$0.6 66.7$\pm$2.4 61.2$\pm$1.1 56.8$\pm$2.1 65.6
DEAL 75.4$\pm$0.3 74.6$\pm$1.1 66.1$\pm$0.6 74.6$\pm$0.8 53.8$\pm$1.0 69.6$\pm$1.8 66.4$\pm$0.3 73.9$\pm$0.6 61.6$\pm$1.4 69.8$\pm$0.2 60.7$\pm$1.0 58.3$\pm$0.9 67.1
CoCo 74.6$\pm$0.9 77.2$\pm$0.6 64.1$\pm$3.4 73.8$\pm$1.1 60.5$\pm$0.2 71.5$\pm$0.7 65.9$\pm$0.5 76.0$\pm$0.5 61.4$\pm$0.4 72.6$\pm$0.6 59.6$\pm$1.0 64.7$\pm$1.0 68.5
SGDA 56.6$\pm$0.6 56.9$\pm$0.8 55.3$\pm$1.2 54.6$\pm$0.5 57.9$\pm$1.3 58.3$\pm$0.4 56.1$\pm$0.9 55.9$\pm$0.6 54.6$\pm$1.3 56.7$\pm$0.5 53.3$\pm$0.7 56.8$\pm$1.1 56.1
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.4$\pm$0.8 56.1$\pm$0.6 56.7$\pm$1.0 55.3$\pm$0.5 54.9$\pm$0.7 57.2$\pm$0.9 55.7$\pm$0.5 56.5$\pm$1.3 54.5$\pm$0.6 56.8$\pm$0.5 56.2$\pm$1.0 58.8$\pm$0.8 56.1
PA-BOTH 56.1$\pm$0.5 56.0$\pm$0.4 56.3$\pm$0.7 56.4$\pm$0.4 56.0$\pm$0.6 57.1$\pm$0.7 56.2$\pm$1.1 58.3$\pm$0.9 56.5$\pm$0.6 57.2$\pm$0.9 56.9$\pm$0.4 57.7$\pm$0.8 56.8
DeSGDA 75.9$\pm$1.0 74.7$\pm$1.2 66.8$\pm$1.5 71.3$\pm$1.9 63.2$\pm$1.5 69.4$\pm$1.7 66.5$\pm$1.3 74.3$\pm$2.4 62.6$\pm$1.5 73.0$\pm$1.8 61.6$\pm$2.2 63.6$\pm$1.5 68.6
[…]
WL subtree 74.9 74.8 67.3 69.9 57.8 57.9 73.7 80.2 60.0 57.9 70.2 73.1 68.1
GCN 73.0$\pm$1.7 68.7$\pm$1.5 66.8$\pm$3.5 69.2$\pm$0.9 53.9$\pm$3.4 53.4$\pm$2.7 69.3$\pm$0.8 74.0$\pm$1.1 55.1$\pm$1.3 42.6$\pm$1.9 55.5$\pm$3.5 57.9$\pm$2.9 61.6
GIN 74.1$\pm$1.8 73.4$\pm$3.4 65.4$\pm$1.5 70.4$\pm$2.9 58.9$\pm$2.7 61.2$\pm$1.1 73.2$\pm$3.8 77.7$\pm$3.0 63.1$\pm$3.7 63.9$\pm$2.4 67.4$\pm$2.3 73.2$\pm$1.9 68.5
GMT 69.0$\pm$4.0 67.4$\pm$3.8 60.3$\pm$4.2 66.5$\pm$3.8 54.9$\pm$1.6 54.8$\pm$3.6 65.6$\pm$4.2 70.4$\pm$3.2 64.0$\pm$2.3 56.8$\pm$4.3 64.7$\pm$1.5 61.1$\pm$3.5 63.0
CIN 68.5$\pm$2.1 65.1$\pm$2.6 65.4$\pm$1.3 63.6$\pm$2.8 57.3$\pm$3.4 59.0$\pm$3.1 59.3$\pm$1.5 68.3$\pm$1.3 58.1$\pm$2.4 71.1$\pm$3.1 60.7$\pm$1.7 61.7$\pm$2.4 63.2
SpikeGCN 66.7$\pm$1.5 65.5$\pm$2.0 57.9$\pm$0.4 60.2$\pm$1.6 53.2$\pm$1.4 60.1$\pm$1.5 57.7$\pm$1.2 67.3$\pm$1.5 57.7$\pm$2.1 60.1$\pm$1.9 59.9$\pm$2.4 63.3$\pm$1.8 60.1
DRSGNN 66.9$\pm$1.2 62.1$\pm$0.7 57.1$\pm$1.2 63.3$\pm$2.1 56.6$\pm$0.9 62.1$\pm$1.3 56.9$\pm$1.0 67.2$\pm$1.8 58.1$\pm$0.6 61.3$\pm$2.5 58.8$\pm$1.0 64.7$\pm$1.7 61.3
CDAN 74.2$\pm$0.3 73.7$\pm$0.5 68.8$\pm$0.2 71.8$\pm$0.4 59.9$\pm$2.0 58.6$\pm$1.9 70.7$\pm$1.4 74.3$\pm$0.3 59.2$\pm$1.2 69.0$\pm$1.6 60.0$\pm$1.2 62.7$\pm$1.3 66.9
ToAlign 75.5$\pm$1.9 67.1$\pm$3.8 68.1$\pm$1.5 63.3$\pm$2.7 55.6$\pm$1.2 67.3$\pm$4.3 69.4$\pm$3.3 77.0$\pm$1.2 57.6$\pm$1.6 74.9$\pm$2.4 59.0$\pm$3.3 64.6$\pm$3.4 66.6
MetaAlign 74.5$\pm$0.9 73.8$\pm$0.6 69.4$\pm$1.2 72.6$\pm$1.3 59.8$\pm$1.8 70.7$\pm$2.7 72.0$\pm$0.5 75.6$\pm$0.6 62.4$\pm$2.1 72.3$\pm$1.9 62.2$\pm$1.1 72.0$\pm$1.2 69.7
DEAL 76.3$\pm$0.2 72.4$\pm$0.7 68.8$\pm$1.0 72.5$\pm$0.7 57.6$\pm$0.6 67.6$\pm$1.9 77.4$\pm$0.6 80.0$\pm$0.7 64.9$\pm$0.7 72.8$\pm$1.4 70.3$\pm$0.3 76.2$\pm$1.3 71.4
CoCo 77.5$\pm$0.4 75.7$\pm$1.3 68.3$\pm$3.7 74.9$\pm$0.5 65.1$\pm$2.1 74.0$\pm$0.4 76.9$\pm$0.6 77.4$\pm$3.4 66.4$\pm$1.5 71.2$\pm$2.7 62.8$\pm$4.2 77.1$\pm$0.6 72.2
SGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
DGDA OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
A2GNN 55.3$\pm$0.3 54.9$\pm$0.6 55.8$\pm$0.4 55.1$\pm$0.8 54.2$\pm$1.0 57.1$\pm$1.2 56.1$\pm$0.5 55.2$\pm$0.7 57.9$\pm$1.5 56.3$\pm$0.6 54.4$\pm$0.5 58.1$\pm$1.5 55.8
PA-BOTH 56.3$\pm$0.5 57.7$\pm$0.9 56.9$\pm$0.6 56.2$\pm$1.0 55.7$\pm$0.8 56.5$\pm$0.9 57.8$\pm$1.2 56.9$\pm$2.1 56.5$\pm$1.5 56.2$\pm$1.8 56.8$\pm$1.4 57.4$\pm$0.7 56.8
DeSGDA 75.8$\pm$1.4 74.5$\pm$1.7 69.5$\pm$1.3 75.0$\pm$2.0 61.0$\pm$1.5 69.2$\pm$1.3 69.5$\pm$1.6 76.1$\pm$1.5 65.0$\pm$1.4 75.5$\pm$2.2 63.4$\pm$1.8 68.3$\pm$1.3 70.3

### v4

@@ Appendix I Baselines
WL subtree 51.5 59.5 28.6 23.8 23.4 19.4 56.8 51.2 54.9 50.2 61.5 57.3 44.8
GCN 49.6$\pm$2.2 62.7$\pm$2.3 22.8$\pm$2.0 26.9$\pm$1.4 13.9$\pm$2.0 22.6$\pm$1.9 74.6$\pm$1.3 58.7$\pm$2.4 75.1$\pm$1.1 52.2$\pm$1.6 76.6$\pm$1.3 67.5$\pm$2.1 50.3
GIN 48.9$\pm$2.8 25.9$\pm$1.8 44.6$\pm$1.5 23.0$\pm$2.0 57.2$\pm$1.8 19.4$\pm$2.0 71.8$\pm$1.2 54.8$\pm$2.4 62.2$\pm$1.5 52.7$\pm$1.9 71.3$\pm$1.7 67.5$\pm$2.4 50.0
GMT 50.8$\pm$2.2 42.7$\pm$2.5 34.9$\pm$2.7 34.8$\pm$1.8 48.2$\pm$2.0 29.6$\pm$2.5 68.9$\pm$1.5 52.6$\pm$1.6 71.2$\pm$1.5 57.1$\pm$2.1 75.9$\pm$1.4 67.8$\pm$1.6 52.9
CIN 50.4$\pm$2.2 29.4$\pm$2.0 23.1$\pm$1.4 31.6$\pm$1.7 42.8$\pm$2.0 24.6$\pm$1.6 54.2$\pm$1.2 57.5$\pm$1.6 73.5$\pm$2.2 52.7$\pm$1.3 75.6$\pm$1.6 67.1$\pm$2.1 48.6
SpikeGCN 56.4$\pm$1.9 70.5$\pm$2.1 34.1$\pm$2.6 53.2$\pm$2.9 20.7$\pm$1.6 49.1$\pm$1.7 79.7$\pm$2.4 66.5$\pm$1.2 77.3$\pm$2.1 61.7$\pm$1.6 78.7$\pm$2.0 71.0$\pm$1.5 59.9
DRSGNN 55.3$\pm$2.4 69.9$\pm$2.2 27.4$\pm$2.0 47.6$\pm$2.7 17.9$\pm$1.6 47.4$\pm$2.1 70.7$\pm$2.0 65.9$\pm$1.7 76.9$\pm$2.1 62.2$\pm$1.4 78.5$\pm$1.8 71.4$\pm$1.6 57.6
CDAN 49.7$\pm$1.9 65.3$\pm$2.3 45.4$\pm$1.8 43.1$\pm$2.1 42.8$\pm$1.7 51.8$\pm$1.6 71.5$\pm$2.0 64.9$\pm$1.6 74.5$\pm$2.5 59.2$\pm$2.2 77.9$\pm$2.1 69.0$\pm$1.5 59.5
ToAlign 52.3$\pm$2.5 66.5$\pm$2.0 47.1$\pm$1.6 45.6$\pm$1.8 41.2$\pm$2.2 51.2$\pm$1.8 73.9$\pm$1.9 65.9$\pm$2.3 77.6$\pm$2.0 60.8$\pm$1.6 78.1$\pm$2.4 70.2$\pm$2.1 60.9
MetaAlign 48.1$\pm$2.0 67.3$\pm$1.7 32.8$\pm$2.0 19.4$\pm$1.8 23.9$\pm$2.5 19.4$\pm$1.7 70.1$\pm$1.8 51.9$\pm$2.1 77.3$\pm$3.2 51.9$\pm$1.6 76.1$\pm$1.8 70.5$\pm$2.0 50.7
DEAL 58.4$\pm$1.5 70.6$\pm$2.0 63.9$\pm$1.6 54.1$\pm$2.1 66.9$\pm$2.4 51.8$\pm$1.6 75.1$\pm$2.5 67.4$\pm$1.8 77.8$\pm$1.9 60.3$\pm$2.1 80.5$\pm$1.8 75.0$\pm$2.0 66.8
CoCo 60.9$\pm$2.3 69.6$\pm$1.2 62.2$\pm$2.2 66.2$\pm$2.0 66.0$\pm$1.8 52.5$\pm$2.3 71.1$\pm$2.4 65.3$\pm$1.5 78.9$\pm$1.3 60.3$\pm$1.4 79.6$\pm$2.1 73.5$\pm$1.8 67.3
SGDA 57.2$\pm$1.5 68.8$\pm$1.8 42.3$\pm$2.0 61.4$\pm$1.7 39.8$\pm$2.2 52.0$\pm$1.8 66.7$\pm$1.9 66.4$\pm$2.3 78.1$\pm$2.1 63.6$\pm$2.6 73.6$\pm$1.6 70.8$\pm$1.9 61.7
StruRW 52.5$\pm$2.5 56.7$\pm$1.3 39.0$\pm$2.3 40.1$\pm$2.0 24.4$\pm$2.1 29.7$\pm$2.4 75.4$\pm$1.7 63.3$\pm$2.0 74.8$\pm$1.6 53.4$\pm$1.5 75.4$\pm$1.4 68.7$\pm$1.7 54.6
A2GNN 53.1$\pm$2.0 65.3$\pm$1.7 42.8$\pm$1.9 40.5$\pm$2.1 33.9$\pm$2.5 39.4$\pm$1.8 69.8$\pm$2.2 61.9$\pm$1.9 77.3$\pm$2.1 61.9$\pm$2.0 76.1$\pm$2.3 67.2$\pm$1.8 57.4
PA-BOTH 51.9$\pm$1.8 50.6$\pm$2.0 35.8$\pm$1.5 37.7$\pm$1.7 27.6$\pm$2.3 43.7$\pm$1.9 62.1$\pm$1.6 61.2$\pm$1.9 65.7$\pm$2.0 58.2$\pm$1.5 73.3$\pm$1.8 69.5$\pm$2.5 53.1
DeSGraDA 62.1$\pm$2.0 72.0$\pm$2.4 69.1$\pm$1.7 71.7$\pm$2.2 68.7$\pm$1.6 58.6$\pm$2.1 76.1$\pm$1.9 66.4$\pm$1.7 80.7$\pm$2.2 63.6$\pm$1.6 81.9$\pm$1.5 78.6$\pm$2.3 70.8
Appendix L Limitation
The proposed DeSGraDA framework, while demonstrating significant improvements in spiking graph domain adaptation, does have some limitations. It assumes a substantial domain shift between the source and target domains, which may not always be applicable in real-world scenarios where domain shifts are minimal. Additionally, the computational complexity introduced by adversarial feature distribution alignment and pseudo-label distillation could become a bottleneck, especially for large-scale datasets. The framework’s sensitivity to hyperparameters, such as time latency and threshold values, also requires careful tuning for different datasets, which may hinder its practical application. Furthermore, while the method provides a generalization bound, its robustness in diverse real-world settings and its ability to address privacy or fairness concerns in sensitive domains remain underexplored. These aspects highlight opportunities for further refinement and broader applicability of DeSGraDA.
Appendix M Broader Impacts
The work on spiking graph domain adaptation offers significant positive societal impacts, particularly in enhancing domain adaptation for graph-based learning tasks in healthcare, biology, and environmental sciences, where it could improve predictions, decision-making, and personalized outcomes. Additionally, its energy-efficient SNNs could contribute to more sustainable AI models. However, potential negative impacts include the risk of reinforcing biases in sensitive applications such as criminal justice or hiring, particularly if the training data is not diverse or representative. The computational complexity may also hinder adoption in resource-limited settings, potentially widening the digital divide. Moreover, privacy concerns in healthcare and finance must be carefully addressed to avoid using sensitive data. In conclusion, while DeSGraDA holds promise, careful attention to fairness, accessibility, and privacy is necessary for its responsible implementation.